# Suppress and Rebalance: Towards Generalized Multi-Modal Face Anti-Spoofing

### 29 Feb 2024 v1

@@ 1 Introduction
Recently, deep learning has been widely applied in FAS, particularly in unimodal scenarios based on the RGB spectrum. While these models perform well in intra-dataset evaluations, their generalizability to unseen attacks and deployment environments remains limited. Despite the improvement of robustness via domain generalization (DG) techniques [19, 67, 48, 47], these approaches still tend to over-fit the attacks presented in source domains.
@@ 2.1 Domain Generalization for FAS
In FAS, domain generalization focuses on training a model using multiple source domains, with the intention that this model can work effectively on unseen target domains [17, 44]. Previous works verify the effectiveness of adversarial training [36, 21, 62], asymmetric triplet loss [19, 28] and controversial learning [47, 37] in developing a shared feature space across domains. Besides, some studies [48, 67, 66] investigate the style information of different domains, aiming to learn domain-invariant features by minimizing distinct style characteristics. Furthermore, meta-learning-based methods [42, 20, 5, 65, 10, 3] are also discussed to simulate domain shifts during training, fostering the learning of a robust representative feature space to such variations. More recent works focus on parameter-efficient transfer learning [6, 17, 46, 7], enabling a pretrained ViT to efficiently adapt to unseen domains with a lower risk of over-fitting. However, these approaches are proposed for unimodal (i.e. RGB) FAS. Due to the ignorance of modality unreliability, there is a gap in adapting them to multi-modal scenarios. Simple multi-modal extensions of these approaches, through early fusion, late fusion, or even the substitution with multi-modal backbones, still exhibit limited performance under multi-modal DG scenarios.
@@ 2.2 Multi-Modal FAS
Multi-modal FAS refers to the use of multiple spectra, e.g., RGB, depth, and infrared, to reveal live and spoofing traces. Due to the complementary information among different modalities, traces that are undetectable in one modality can be easily captured in others [35]. To fully exploit information from each modality, early methods employ channel-wise concatenation of inputted modalities [14, 40], or develop multiple separate branches for modality feature extraction followed by late fusion [56, 45, 24, 25]. Recent works introduce attention-based feature fusion techniques [26, 4] and adaptive cross-modal loss functions [13] to encourage the extraction of complementary information among modalities. Meanwhile, cross-modality translation is discussed [34, 27] to mitigate the semantic gap between different modalities. To further enhance the practical usability of FAS, Yu et al. [60, 58, 59] introduce flexible-modal benchmarks that force the use of incomplete modalities during training or testing. In the context of flexible-modal FAS, cross-modality attention [35, 32] or multi-modal adapters [59, 58] are proposed for pretrained ViT to learn modality-agnostic live/spoofing features. However, the aforementioned methods fail to perform well in DG scenarios. The lack of sensitivity to modality unreliability and imbalance leads to an insufficient ability to resist domain shifts. Although previous works rebalance modalities by modulating gradients [41] or adjusting modality losses [11], these methods ignore domain shifts and are not suitable for FAS frameworks with complex cross-modal fusion modules.
@@ 3.1 Fine-tuning ViT with U-Adapters
In multi-modal FAS, DG scenarios involving significant domain shifts [31], including unseen face presentations attacks [49], different kinds of imaging sensors, and noise from low-quality sensors [63], often lead to the extraction of unreliable live/spoofing features from certain modal inputs. These unreliable features can propagate during cross-modal fusion, adversely affecting other modalities. To this end, we aim to identify unreliable local features within each modality and suppress the interaction of these features with other modalities, ensuring that each modality utilizes reliable information from other modalities as complementary traces. Specifically, we adopt uncertainty estimated via Monte Carlo sampling [12, 22, 38] to evaluate the feature unreliability in each modality, and based on the uncertainty, we reduce their weight in cross-modal attention-based feature fusion, minimizing their negative impact. Inspired by recent success in parameter-efficient transfer learning (PETL) by introducing adapters into pretrained ViT [17, 6, 7, 59, 58], but distinctively, we propose the U-Adapter for fine-tuning, which enables cross-modal feature fusion and recognizes as well as suppresses the propagation of unreliable information from uncertainly detected regions caused by domain shifts. In Fig. 3, we present the process of fine-tuning ViT with the U-Adapters, including adopting an Uncertainty Estimation Module (UEM) to recognize unreliable tokens and employing the U-Adapter for uncertainty-aware cross-modal fusion. The ViT [9], comprising $n$ transformer blocks each with Layer Normalization (LN), Multi-Head Self Attention (MHSA), and Multi-Layer Perceptron (MLP), is fine-tuned at the MLP’s output using our U-Adapter. By representing our U-Adapter as $\mathcal{A}$ and taking RGB (R), depth (D), and infrared (I) as the inputted modalities, the fine-tuning process is formulated as follows.

### 05 Mar 2024 v2

@@ 1 Introduction
Recently, deep learning has been widely applied in FAS, particularly in unimodal scenarios based on the RGB spectrum. While these models perform well in intra-dataset evaluations, their generalizability to unseen attacks and deployment environments remains limited [2]. Despite the improvement of robustness via domain generalization (DG) techniques [20, 70, 50, 49, 28], these approaches still tend to over-fit the attacks presented in source domains.
@@ 2.1 Domain Generalization for FAS
In FAS, domain generalization focuses on training a model using multiple source domains, with the intention that this model can work effectively on unseen target domains [18, 46]. Previous works verify the effectiveness of adversarial training [37, 22, 65], asymmetric triplet loss [20, 30] and controversial learning [49, 38] in developing a shared feature space across domains. Besides, some studies [50, 70, 69] investigate the style information of different domains, aiming to learn domain-invariant features by minimizing distinct style characteristics. Furthermore, meta-learning-based methods [44, 21, 4, 68, 10] are also discussed to simulate domain shifts during training, fostering the learning of a robust representative feature space to such variations. More recent works focus on parameter-efficient transfer learning [5, 18, 48, 6], enabling a pretrained ViT to efficiently adapt to unseen domains with a lower risk of over-fitting. However, these approaches are proposed for unimodal (i.e. RGB) FAS. Due to the ignorance of modality unreliability, there is a gap in adapting them to multi-modal scenarios. Simple multi-modal extensions of these approaches, through early fusion, late fusion, or even the substitution with multi-modal backbones, still exhibit limited performance under multi-modal DG scenarios.
@@ 2.2 Multi-Modal FAS
Multi-modal FAS refers to the use of multiple spectra, e.g., RGB, depth, and infrared, to reveal live and spoofing traces. Due to the complementary information among different modalities, traces that are undetectable in one modality can be easily captured in others [36, 17]. To fully exploit information from each modality, early methods employ channel-wise concatenation of inputted modalities [14, 42], or develop multiple separate branches for modality feature extraction followed by late fusion [58, 47, 25, 26]. Recent works introduce attention-based feature fusion techniques [27, 8] and adaptive cross-modal loss functions [13] to encourage the extraction of complementary information among modalities. Meanwhile, cross-modality translation is discussed [35, 29] to mitigate the semantic gap between different modalities. To further enhance the practical usability of FAS, Yu et al. [62, 60, 61] introduce flexible-modal benchmarks that force the use of incomplete modalities during training or testing. In the context of flexible-modal FAS, cross-modality attention [36, 33] or multi-modal adapters [61, 60] are proposed for pretrained ViT to learn modality-agnostic live/spoofing features. However, the aforementioned methods fail to perform well in DG scenarios. The lack of sensitivity to modality unreliability and imbalance leads to an insufficient ability to resist domain shifts. Although previous works rebalance modalities by modulating gradients [43] or adjusting modality losses [11], these methods ignore domain shifts and are not suitable for FAS frameworks with complex cross-modal fusion modules.
@@ 3.1 Fine-tuning ViT with U-Adapters
In multi-modal FAS, DG scenarios involving significant domain shifts [32], including unseen face presentations attacks [51], different kinds of imaging sensors, and noise from low-quality sensors [66], often lead to the extraction of unreliable live/spoofing features from certain modal inputs. These unreliable features can propagate during cross-modal fusion, adversely affecting other modalities. To this end, we aim to identify unreliable local features within each modality and suppress the interaction of these features with other modalities, ensuring that each modality utilizes reliable information from other modalities as complementary traces. Specifically, we adopt uncertainty estimated via Monte Carlo sampling [12, 23, 40] to evaluate the feature unreliability in each modality, and based on the uncertainty, we reduce their weight in cross-modal attention-based feature fusion, minimizing their negative impact. Inspired by recent success in parameter-efficient transfer learning (PETL) by introducing adapters into pretrained ViT [18, 5, 6, 61, 60, 39], but distinctively, we propose the U-Adapter for fine-tuning, which enables cross-modal feature fusion and recognizes as well as suppresses the propagation of unreliable information from uncertainly detected regions caused by domain shifts. In Fig. 3, we present the process of fine-tuning ViT with the U-Adapters, including adopting an Uncertainty Estimation Module (UEM) to recognize unreliable tokens and employing the U-Adapter for uncertainty-aware cross-modal fusion. The ViT [9], comprising $n$ transformer blocks each with Layer Normalization (LN), Multi-Head Self Attention (MHSA), and Multi-Layer Perceptron (MLP), is fine-tuned at the MLP’s output using our U-Adapter. By representing our U-Adapter as $\mathcal{A}$ and taking RGB (R), depth (D), and infrared (I) as the inputted modalities, the fine-tuning process is formulated as follows.